OpenStudy (anonymous):

Trella Alcala deposited 1,950.00 in a new credit union savings account on the first of the quarter. The principal earns 4.25% interest compounded quarterly. She made no other deposits or withdrawals. What was the amount in her account at the end of 6 months? What was the compound interest? Please explain to me how to get the answer.

OpenStudy (anonymous):

6 month is equal to 2 quarters

OpenStudy (anonymous):

so A=1950(0.045)^2

OpenStudy (anonymous):

1950(1.045)^2

ganeshie8 (ganeshie8):

P = 1,950 r = 4.25 % = .0425 n = 6 months = 2 quarters t = 1/2 years use the formula, \(\huge Pnew = P(1+\frac{r}{n})^{n*t} \)

ganeshie8 (ganeshie8):

oops - typo there n = 4 months = 4

OpenStudy (anonymous):

oh ok, Thank you. I did not understand what quarterly was. That was really confusing me.

ganeshie8 (ganeshie8):

\(\huge Pnew = 1950 * (1+\frac{.0425}{4})^{4*\frac{1}{2}}\)

ganeshie8 (ganeshie8):

np :)

OpenStudy (anonymous):

I will try out your formula. Thank you.